Draw the graphs of 3x+7=0 and 2y-3=0. Do these lines intersect? What relationship do you observe between these lines?

Answer by stanbon(75887)

Draw the graphs of 3x+7=0 and 2y-3=0. Do these lines intersect? What relationship do you observe between these lines?
----
x = -7/3
y = 3/2
-----
lines are perpendicular
------------------------
Cheers,
Stan H.
--------------